OpenStudy (anonymous):

A biologist is comparing the growth of a population of flies per week to the number of flies a lizard will consume per week. She has devised an equation to solve for which day (x) the lizard would be able to eat the entire population. The equation is 3^x = 5x − 1. Explain to the biologist how she can solve this on a graph using a system of equations.

OpenStudy (anonymous):

@whpalmer4, anyone?

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@prowrestler

OpenStudy (anonymous):

@Michele_Laino please please help me or ask someone for me

OpenStudy (michele_laino):

A possible way is: I call with g(x) the function 3^x, and with g(x) the function 5x-1, namely I write this: \[\Large \begin{gathered} f\left( x \right) = {3^x} \hfill \\ g\left( x \right) = 5x - 1 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

then I draw the graph of both functions f(x) and g(x) and I search for intersection point of those graphs

OpenStudy (anonymous):

is there a way of finding the intersection without graphing

OpenStudy (michele_laino):

one point is given setting x=2 we have: \[\Large \begin{gathered} f\left( 2 \right) = {3^2} = 9 \hfill \\ g\left( 2 \right) = 5 \cdot 2 - 1 = 9 \hfill \\ \end{gathered} \] so the corresponding intersection point is: \[\Large \left( {2,9} \right)\]

OpenStudy (michele_laino):

another point can be compute, if we expand the function f(x) around x=0, using Taylor expansion

OpenStudy (michele_laino):

computed*

OpenStudy (anonymous):

where did you get 2 from

OpenStudy (michele_laino):

I did some trial

OpenStudy (anonymous):

oh ok and thank you

OpenStudy (michele_laino):

please wait, try to write the Taylor expansion, around, x=0 of f(x), or try to use a software online like "desmos"

OpenStudy (anonymous):

hold the question says use system of equations

OpenStudy (michele_laino):

yes! in fact I broke your equation in two functions

OpenStudy (michele_laino):

here is the system: \[\Large \left\{ \begin{gathered} f\left( x \right) = {3^x} \hfill \\ g\left( x \right) = 5x - 1 \hfill \\ \end{gathered} \right.\]

OpenStudy (anonymous):

but there is no way of eliminating anything so is that why we graph

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it is just that i did not learn about the taylor expansion

OpenStudy (michele_laino):

please here is the expansion up to the first order term:

OpenStudy (michele_laino):

\[\Large {3^x} \simeq {\left. {{3^x}} \right|_{x = 0}} + {\left. {{3^x}\log 3} \right|_{x = 0}}x = 1 + x\log 3\]

OpenStudy (michele_laino):

now, if we want to solve your equation, namely: 3^x=5x-1, we can solve this equation: \[\Large 1 + x\log 3 = 5x - 1\]

OpenStudy (michele_laino):

so we get the second intersection point

OpenStudy (anonymous):

ok thank you, no need to go further.

OpenStudy (michele_laino):

:)